# Joint Beamforming and Position Optimization for Fluid RIS-aided ISAC Systems

### v2

@@ III-F Convergence and Computational Complexity Analysis
The proposed algorithm is summarized in Algorithm 2, which incorporates the solutions derived above. The algorithm iteratively optimizes the variables, where the process continues until the objective function value in (19) becomes stable.
@@ III-F1 Convergence Analysis
In this part, we analyze that the convergence of the proposed Algorithm 2. First, the subproblem (20) is a unconstrained convex quadratic problem with a close-form solution, and thus $\omega^{t}$ update does not increase objective function value:
$\varepsilon(\omega^{t+1},\mathbf{\Theta}^{t},\mathbf{W}^{t},\mathbf{p}^{t})\leq\varepsilon(\omega^{t},\mathbf{\Theta}^{t},\mathbf{W}^{t},\mathbf{p}^{t}).$ (45)
Updating $\mathbf{\Theta}^{t}$ in (23), no matter utilizing the SDR [39] or the manifold method [40], it was proved that the objective function value of (23) is non-increasing over iterations, such that
$\varepsilon(\omega^{t+1},\mathbf{\Theta}^{t+1},\mathbf{W}^{t},\mathbf{p}^{t})\leq\varepsilon(\omega^{t+1},\mathbf{\Theta}^{t},\mathbf{W}^{t},\mathbf{p}^{t}).$ (46)
In the optimization of $\mathbf{W}$, quasi-Newton method guarantees that the update of $\overline{\mathbf{w}}$ is non-increasing [42], while equality constraint is satisfied if the penalty parameter becomes sufficiently large. As a result, the solution to (33) can satisfy the equality constraint, while the process (33) is non-increasing:
$\varepsilon(\omega^{t+1}\!,\mathbf{\Theta}^{t+1}\!,\mathbf{W}^{t+1}\!,\mathbf{p}^{t})\!\leq\!\varepsilon(\omega^{t+1}\!,\mathbf{\Theta}^{t+1}\!,\mathbf{W}^{t}\!,\mathbf{p}^{t}).$ (47)
Proved in [25], the subproblem (44) is a non-decreasing process given the nature of its convexity and the inequalities of Taylor expansions, while the constrained optimization process can converge in the MM framework when the function is smooth objective function [47]:
$\varepsilon(\omega^{t+1}\!,\mathbf{\Theta}^{t+1}\!,\mathbf{W}^{t+1}\!,\mathbf{p}^{t+1})\!\leq\!\varepsilon(\omega^{t+1}\!,\mathbf{\Theta}^{t+1}\!,\mathbf{W}^{t+1}\!,\mathbf{p}^{t}).$ (48)
Besides, the finite resources indicate that the objective function value is bounded and greater than zero. As a result, the Algorithm 2 can guarantee to converge.
@@ III-F2 Computational Complexity Analysis
  • To obtain $\mathbf{s}_{r}$, the computational complexity mainly stems from solving the problem (16) with SDR and the eigenvalue decomposition. Given a solution accuracy $\epsilon_{1}$, the computational complexity of SDR is $\mathcal{O}\left(N^{4.5}\log(1/\epsilon_{1})\right)$ [36], while the eigenvalue decomposition has the computational complexity of $\mathcal{O}\left(N^{3}\right)$.
  • The optimal symbol estimator can be calculated via a close-form expression, whose computational complexity stems from matrix multiplications. Obtaining the numerator in (21) has the computation complexity of $\mathcal{O}\left(KN+N^{2}+MN+M\right)$, while calculating the denominator has the computation complexity of $\mathcal{O}\left(2(MK+NK+N^{2})+M\right)$.
  • When optimize the fRIS phase shift, the computational complexity of calculation of the matrix multiplication is $\mathcal{O}\left(MN+MK+N^{2}K+N^{2}M+MNK\right)$. Besides, the SDR approach has the computational complexity of $\mathcal{O}\left(N^{4.5}\log(1/\epsilon_{2})\right)$ with $\epsilon_{2}$ being the solution accuracy, while that of the manifold optimization is $\mathcal{O}\left(N^{1.5}\right)$ [40].
  • In the optimization of transmit beamformer design, the computational complexity of matrix multiplication is $\mathcal{O}\left(M^{2}K^{2}\!+\!M^{2}N\!+\!M^{2}K\!+\!N^{2}M\!+\!MK\!+\!MN\!+\!K^{2}\right)$, while the design of the transmit beamformer with ALM is $\mathcal{O}\left((M^{2}+M)\log\left(1/\epsilon_{3}\right)\right)$ with $\epsilon_{3}$ being the convergence accuracy [48, 49].
  • When optimize the position of one fRIS element, the computational complexity of obtaining the gradients and the Hessian matrix is $\mathcal{O}\left(NK+K+M+N^{2}\right)$, while the computational complexity of calculating $\delta_{n}$ is $\mathcal{O}\left(1\right)$. Additionally, the computational complexity of solving the problem (44) is $\mathcal{O}\left(\left(N^{1.5}\log(1/\epsilon_{4})\right)\right)$ [25] with accuracy $\epsilon_{4}$ for the interior-point method.
As a result, the total computational complexity of each iteration of the Algorithm 2 can be approximated as $\mathcal{O}\left(N^{4.5}\log(1/\epsilon_{1})\right)$, neglecting the lower-order terms.
@@
0:    Channel parameters, noise power, transmit power budget, ideal sensing beampattern, weighting coefficient, users number, fRIS planar region size, and the minimum inter-element spacing requirement.
1:  Initialize all the variables;
2:  while no convergence of $\varepsilon$ do
3:     Calculate correlation matrix for sensing-only by (16) and obtain desired reflective signal $\mathbf{s}_{r}$ via (17);
@@
6:     Update ${\mathbf{W}}$ by implementing the ALM process in (33);
@@
8:  end while
9:  return  Optimized variables $\mathbf{W}$, $\boldsymbol{\Theta}$, $\mathbf{p}$, $\omega$, $\beta$ and the objective function value.
@@ IV Simulation Results
This section provides numerical results to validate the effectiveness of the fRIS-aided ISAC system and the proposed algorithm. Unless stated otherwise, the simulation parameters are set as follows. The center frequency of this system is $2.4$GHz, corresponding to a wavelength of $\lambda\!=\!0.125$m. The ISAC BS is located at ($3$m, 00m, 00m), and equipped with $M\!=\!8$ antennas. The transmitter has a maximum transmit power of $10$dBm. The fRIS, consisting of $16$ elements ($N_{x}\!=\!N_{y}=4$), is positioned at (00m, $3$m, $3$m), with a size of $A\!\times\!A$ where $A\!=\!4\lambda$ and a minimal inter-element spacing of $\Delta D=\lambda/2$. Additionally, we consider $K\!\!=\!\!4$ communication users and $T\!\!=\!\!3$ targets. The users are randomly distributed within an area centered at ($30$m, $100$m, 00m) with a radius of $10$m. As observed from fRIS, the targets’ (azimuth, elevate) angles are ($-60^{\circ}$, $0^{\circ}$), ($10^{\circ}$, $0^{\circ}$) and ($55^{\circ}$, $0^{\circ}$), respectively. The noise power of the users are set as $\sigma_{0}^{2}\!=\!-60$dBm and QPSK is used for modulation. The weighting coefficient $\alpha$ for balancing the communication and sensing performance is $0.5$. Moreover, the channel fading coefficients $\zeta_{G}$ and $\zeta_{k}$ are calculated by $\eta\cdot(\frac{1}{\text{dist}})^{2}$, where $\eta\!=\!-10$dB is the path loss exponent and dist denotes the transmitter-receiver distance [39]. The penalty factor for ALM and the convergence threshold are set as $1$ and $10^{-5}$, respectively. All simulations are averaged over $100$ Monte Carlo trials.
Figure 2: Convergence behaviors of the proposed algorithm and the benchmarks: the objective function value $\varepsilon$ versus the iteration step.
Figure 3: The effect of RIS element number: the objective function value $\varepsilon$ versus the element number.
@@
  • Conventional RIS-aided ISAC system (Conven): The positions of RIS elements are fixed and uniformly distributed across the fRIS plane. The rest of the beamforming designs follow the same procedures as proposed.
  • Discrete position selection (DPS): The moving region of the elements is quantized into discrete grids of equal spacing, where greedy algorithm is employed for the position selection. Specifically, the position of each element is selected sequentially in the discrete position set, while the positions of the other elements are kept unchanged.
  • Random RIS element deployment (Rand): The positions of RIS elements remain fixed for each channel realization but randomly vary across realizations. The optimization of other components is performed as the proposed scheme.
@@
Figure 4: Beampattern at the fRIS: beam gain (dB) versus spatial angle (degree). (a) The beampattern when bias to the communication ($\alpha=0.1$). (b) The beampattern when no bias ($\alpha=0.5$). (c) The beampattern when bias to the sensing ($\alpha=0.9$). (d) The beampattern of different approaches when $\alpha=0.5$.
TABLE I: Comparisons in terms of Average Performance Value
and Time Complexity per round
Scheme Objective Function Value Elapsed Time (s)
Proposed 0.378 313.52
($N=25$)
Conven 0.394 934.01
($N=64$)
Conven 0.432 372.46
($N=49$)
In Fig.2, we examine the convergence performance of our proposed algorithm for the fRIS-assisted ISAC system and compare it with the benchmark scenarios. As expected, the proposed approach demonstrates convergence, with the objective function value decreasing rapidly and stabilizing at $0.57$ after $15$ iterations. Meanwhile, it can be observed that the joint MSE of the proposed fRIS-aided ISAC case outperforms the baselines, especially compared to the conventional RIS-aided ISAC case. This improvement can be attributed to the continuous and flexible element position optimization, which provides additional DoFs in the spatial domain.
@@
Fig.3 illustrates the relationship between the number of fRIS elements ($N$) and the objective function value ($\varepsilon$), with $N$ varying from $16$ to $64$. As observed, the objective function value of the proposed scheme and the baselines all decrease as $N$ increases, while the proposed scheme achieves the best performance among all. In addition, in Table I, we compare the average objective function value $\varepsilon$ and the time complexity of the proposed scheme with those of the conventional RIS-aided ISAC system. Specifically, we select the proposed scheme with $N=25$, the “Conven” scheme with $N=49$ and $N=64$, where either the objective function value or the elapsed time has similar performance. When compared with the “Conven” scheme with $N=64$, the proposed scheme with $N=25$ can achieve slightly better objective function value yet requiring $1/3$ of the elapsed time. When compared with the “Conven” scheme with $N=49$, the proposed scheme with $N=25$ exhibits comparable elapsed time but significantly better objective function value performance. These results demonstrate that employing the fRIS not only compensates for performance losses caused by reducing RIS elements but also alleviates the complexity of high-dimensional optimization.
@@
To evaluate the system’s sensing performance, we further investigate the fRIS beampattern under different weighting coefficients, and compare with the ideal case and the baselines in Fig.4. It is worth noting that the targets are assumed to lie at an angle of $0^{\circ}$ in elevate, so that the beampatterns of the cross section at $0^{\circ}$ in elevate are plotted for clearer observations and easier comparisons. In Fig.4 - Fig.4, the practical, sensing-only and ideal beampatterns are illustrated for the weighted coefficients $\alpha=0.1,0.5\text{ and }0.9$, respectively. It can be observed that beams are consistently formed in the target directions across all $\alpha$ values. However, in the case of $\alpha=0.1$ where the system is biased towards communication, the beam is not as similar as the sensing-only case and exhibits higher sidelobe levels. In contrast, when $\alpha=0.5$ and $\alpha=0.9$ where the system is increasingly biased towards the sensing, the practical beams gradually align with the sensing-only case. On the other hand, Fig.4 compares the beampattern obtained from different approaches when $\alpha=0.5$. It can be found that all the beampatterns have formed peaks in the target directions, but the proposed method achieves lower sidelobe levels compared to the conventional case and higher peaks compared to the ”DPS” and ”Rand” cases.
@@
Figure 5: Average BER (dB) versus different receive noise power (dBm) under different modulations and weighting coefficients.
Figure 6: Average BER (dB) versus different receive noise power (dBm) with different approaches.
@@
In addition, we evaluate the communication performance of the system under different settings. Fig.5 depicts the average bit error rate (BER) of the proposed scheme versus various noise power levels, ranging from $-40$dBm to $-80$dBm. As expected, the average BER decreases when the noise power at the receivers decreases, and QPSK modulation exhibits lower BER when compared to 16QAM. Moreover, when $\alpha$ decreases, the BER further improves as the system becomes more communication-biased. Notably, at a noise power of $-80$dBm and a weighting coefficient of $\alpha=0.9$, the BER can drop to approximately $-28$dB with QPSK modulation. On the other hand, Fig. 6 compares the average BER of the proposed method to the benchmarks. As shown, all cases exhibit a decreasing BER trend as noise power level reduces, while the proposed method consistently achieves a lower BER compared to the baselines.
Figure 7: The effect of the fRIS region size: the objective value versus the normalized region size $A/\lambda$.
In Fig.7, the impact of the normalized region size $A/\lambda$ on the proposed scheme is illustrated, where $A$ varies from $2\lambda$ to $6\lambda$. The results show that the objective function value $\varepsilon$ decreases as the region size increases, which is expected since the larger size of the fRIS provides more DoFs for the mobility of the elements. On the other hand, the benchmarks of “DPS” and “Rand” cases demonstrate a decreasing trend with increasing region size, while the “Conven” case has little change as the region size varies. Compared with the benchmarks, the proposed scheme and the optimization algorithm achieve superior performance across all region sizes, which is consistent to the previous simulation results.

### v3

@@ III-F Convergence and Computational Complexity Analysis
The proposed algorithm is summarized in Algorithm 2, which will terminate when the change of $\varepsilon$ falls below a threshold $\epsilon_{0}$ or the maximum iteration number $I_{\mathrm{max}}$ is reached.
@@ III-F1 Convergence Analysis
The convergence behavior of the proposed Algorithm 2 is analyzed as below.
(i) Subproblem of symbol estimator optimization: The update of $\omega$ in (19) solves a convex quadratic problem, which guarantees $\varepsilon(\omega^{t+1},\mathbf{\Theta}^{t},\mathbf{x}^{t},\mathbf{p}^{t})\leq\varepsilon(\omega^{t},\mathbf{\Theta}^{t},\mathbf{x}^{t},\mathbf{p}^{t})$.
(ii) Subproblem of phase shift optimization: The subproblem (22) involves the modulus-one constraint. It is solved via either SDR [42] or manifold optimization [49], where the SDR is emphatically shown as a non-increasing process and manifold optimization updates variables along the Riemannian gradient direction on the complex circle manifold. Both approaches ensure $\varepsilon(\omega^{t+1},\mathbf{\Theta}^{t+1},\mathbf{x}^{t},\mathbf{p}^{t})\leq\varepsilon(\omega^{t+1},\mathbf{\Theta}^{t},\mathbf{x}^{t},\mathbf{p}^{t})$.
(iii) Subproblem of transmit waveform optimization: In the optimization of $\mathbf{x}$, the quasi-Newton method guarantees that the update of $\overline{\mathbf{x}}$ is non-increasing [3], while the equality constraint is satisfied if the penalty parameter becomes sufficiently large. As a result, the solution to (30) can satisfy the equality constraint, while the process is non-increasing: $\varepsilon(\omega^{t+1}\!,\mathbf{\Theta}^{t+1}\!,\mathbf{x}^{t+1}\!,\mathbf{p}^{t})\!\leq\!\varepsilon(\omega^{t+1}\!,\mathbf{\Theta}^{t+1}\!,\mathbf{x}^{t}\!,\mathbf{p}^{t})$.
(iv) Subproblem of element position optimization: The update of $\mathbf{p}$ is tackled via the MM framework, where a convex quadratic surrogate function is constructed. According to the convergence theory of MM methods [34], the surrogate ensures that the original objective is non-increasing, i.e., $\varepsilon(\mathbf{p}_{n}^{(i-1)})=g(\mathbf{p}_{n}^{(i-1)})+C\geq g(\mathbf{p}_{n}^{(i)})+C=\varepsilon(\mathbf{p}_{n}^{(i)})$, where $g(\cdot)$ is the surrogate function while $C$ is a constant. Besides, the inter-element spacing constraint is relaxed to a half space, which does not affect the monotonicity of the descent.
Moreover, due to the non-negativity of MSEs and the finite transmit power, the overall objective function is lower-bounded by 0. Hence, the proposed Algorithm 2 generates a non-increasing and bounded sequence of objective values, and converges to a stationary point that is at least a local optimum.
@@ III-F2 Computational Complexity Analysis
  • To obtain $\mathbf{s}_{r}$, the computational complexity mainly stems from solving (15) with SDR and the eigenvalue decomposition. Given a solution accuracy $\epsilon_{1}$, the computational complexity of SDR is $\mathcal{O}\left(N^{4.5}\log(1/\epsilon_{1})\right)$ [21], while the eigenvalue decomposition has the computational complexity of $\mathcal{O}\left(N^{3}\right)$.
  • The optimal symbol estimator can be calculated via a closed-form expression, whose computational complexity stems from matrix multiplications. Obtaining the numerator in (20) has the computation complexity of $\mathcal{O}\left(KN+N^{2}+MN+M\right)$, while calculating the denominator has the computation complexity of $\mathcal{O}\left(2(MK+NK+N^{2})+M\right)$.
  • When optimizing the fRIS phase shift, the computational complexity of the calculation of the matrix multiplication is $\mathcal{O}\left(MN+MK+N^{2}K+N^{2}M+MNK\right)$. Besides, the SDR approach has the computational complexity of $\mathcal{O}\left(N^{4.5}\log(1/\epsilon_{2})\right)$ with $\epsilon_{2}$ being the solution accuracy, while that of the manifold optimization is $\mathcal{O}\left(N^{1.5}\right)$ [49].
  • In the optimization of transmit waveform design, the computational complexity of matrix multiplication is $\mathcal{O}\left(\!M^{2}N\!+\!M^{2}K\!+\!N^{2}M\!+\!MK\!+\!MN\!\right)$, while the design of the transmit waveform with ALM is $\mathcal{O}\left((M^{2}+M)\log\left(1/\epsilon_{3}\right)\right)$, where $\epsilon_{3}$ is the convergence accuracy [14, 11].
  • When optimizing the position of one fRIS element, the computational complexity of obtaining the gradients and the Hessian matrix is $\mathcal{O}\left(NK+K+M+N^{2}\right)$, while the computational complexity of calculating $\delta_{n}$ is $\mathcal{O}\left(1\right)$. Additionally, the computational complexity of solving the problem (42) is $\mathcal{O}\left(\left(N^{1.5}\log(1/\epsilon_{4})\right)\right)$ [23] with accuracy $\epsilon_{4}$ for the interior-point method.
In summary, the computational complexity of each iteration of the Algorithm 2 can be approximated as $\mathcal{O}\left(N^{4.5}\log(1/\epsilon)\right)$, neglecting the lower-order terms.
@@
0:   Channel parameters, noise power, transmit power, ideal sensing beampattern, weighting coefficient, fRIS size, and the minimum inter-element spacing.
1: Initialize all the variables, and $i=0$;
2:while ${|\varepsilon^{(i+1)}-\varepsilon^{(i)}|}/{\varepsilon^{(i)}}>\epsilon_{0}$ or the maximum iteration number is not reached, do
3:  Calculate correlation matrix by (15) and obtain reference signal $\mathbf{s}_{r}$ via (16);
@@
6:  Update $\mathbf{x}$} by implementing the ALM process in (30);
@@
8:  Update iteration index $i=i+1$;
9:end while
10:return Optimized variables and joint MSE.
@@ IV Simulation Results
This section provides numerical results to validate the effectiveness of the system and the proposed algorithm. We set the default parameters as follows. The center frequency of this system is 2.42.4GHz. The ISAC BS is located at ($3$, 0, 0)m, and equipped with M=8M\!=\!8 antennas. The transmitter has maximum transmit power of 1010dBm. The fRIS, consisting of $25$ elements ($N_{x}\!=\!N_{y}=5$), is positioned at (0, 33, 33)m, with a size of $A\!\times\!A$ where $A\!=\!5\lambda$ and a minimal inter-element spacing of $\Delta D=\lambda/2$. Additionally, we consider K=4K\!\!=\!\!4 users and T=3T\!\!=\!\!3 targets. The users are randomly distributed within an area centered at ($50$, $100$, 0)m with a radius of 1010m. Observed from fRIS, the targets’ (azimuth, elevation) angles are ($20^{\circ}$, $0^{\circ}$), ($5^{\circ}$, $0^{\circ}$), and ($30^{\circ}$, $0^{\circ}$), respectively. The noise power is set as $\sigma_{0}^{2}\!=\!-60$dBm, and QPSK is used for modulation. The weighting coefficient $\alpha$ is $0.5$. Moreover, the channel fading coefficients $\zeta_{G}$ and $\zeta_{k}$ are calculated by $\eta\cdot(\frac{1}{\text{dist}})^{2}$, where $\eta\!=\!-3$dB is the path loss, and dist denotes the transmitter-receiver distance [42]. The penalty factor for ALM and the convergence threshold are set as $1$ and $10^{-5}$, respectively. All simulation results are averaged over $100$ Monte Carlo trials. In addition, $\mathbf{x}$, $\boldsymbol{\Theta}$, and $\omega$ are randomly initialized under their respective constraints, while the fRIS element positions $\mathbf{p}$ are initialized as stated in Section III-E.
(a)                                     (b)
Figure 2: Convergence behaviors: (a) The proposed algorithm and the benchmarks: the objective function value $\varepsilon$ versus the iteration step. (b) Fluctuations with a fixed channel realization and random initialization parameters.
@@
  • Conven (SameNum): This scheme is an ISAC aided by fixed-element RIS, where the RIS has the same element number as fRIS (sparse). The rest of the beamforming designs follow the same procedures as proposed.
  • Conven (HWL): The ISAC system is aided by a fixed-element RIS, where the elements have dense half-wavelength spacing in the same surface size. The rest of the optimization procedures are identical to the proposed.
  • DPS: In this scheme, the fRIS element position is discrete, where the fRIS is quantized into discrete grids of equal spacing. Specifically, the continuous position is projected to the nearest position.
@@
Figure 3: Beampattern at the fRIS: beam gain (dB) versus spatial angle (). (a) The beampattern when biased to the communication ($\alpha=0.1$). (b) The beampattern when no bias ($\alpha=0.5$). (c) The beampattern when biased to the sensing ($\alpha=0.9$). (d) The beampattern of different approaches when $\alpha=0.5$.
Figure 4: Beampatterns (dB) of different system configurations: (a) Angles Separation: $5^{\circ}$, Element number: $16$, Size: $8\lambda$. (b) Angles Separation: $10^{\circ}$, Element number: $16$, Size: $8\lambda$. (c) Angles Separation: $5^{\circ}$, Element number: $16$, Size: $12\lambda$. (d) Angles Separation: $5^{\circ}$, Element number: $64$, Size: $8\lambda$. (e) Angles Separation: $10^{\circ}$, Element number: $64$, Size: $8\lambda$. (f) Angles Separation: $5^{\circ}$, Element number: $64$, Size: $12\lambda$.
TABLE I: Confirmation of the joint effect of position and phase optimization of fRIS
Objective function Commun. Sensing
Value MSE MSE
Joint 0.27 0.23 0.31
optimization
Position 0.79 1.1 0.48
only
Phase shift 0.57 0.67 0.47
only
No phase shift or 2.07 3.51 0.63
position optimization
In Fig. 2(a), we examine the convergence performance of the proposed algorithm for the fRIS-aided ISAC system and compare it with the benchmarks. As expected, the proposed approach demonstrates rapid convergence, stabilizing after $20$ iterations. Besides, the MSE of the proposed scheme outperforms the baselines under the same element number, while yielding comparable performance as the “Conven (HWL)” case. This is because the continuous and flexible element position optimization provides additional spatial DoFs. In addition, the DoFs of element position optimization are validated in Table I. As compared to position-only/phase-only optimization, the joint optimization outperforms the individual optimization schemes, implying that repositioning does not overlap with the existing optimization dimension. To assess the impact of initialization, we conduct $100$ trials under a fixed channel realization and different initialization. As shown in Fig. 2(b), the convergence exhibits small variation, with a mean of $0.1461$, a variance of $0.0065$, and a range of [$0.0577$, $0.4922$]. This indicates good robustness and stable convergence behavior.
TABLE II: Comparisons in terms of Average Performance Value
and Time Complexity per iteration
Scheme Objective Function Value Elapsed Time
(s/iteration)
Proposed: 0.27 24.6
$N=25$
Conven (SameNum): 0.57 1.33
$N=25$
Conven (HWL): 0.32 47.75
$N=81$
@@
To further validate the advantages of fRIS, we compare the average objective function value and the time complexity of the proposed scheme with those of the conventional RIS-aided schemes in Table II. Specifically, in the “Conven (SameNum)” scheme, the conventional RIS has the same element number as fRIS ($N=25$), while in the “Conven (HWL)” scheme, the elements have fixed half-wavelength spacing under the same aperture, where the element number is $N=81$. It is seen that the proposed scheme achieves better results compared to “Conven (SameNum)” due to additional DoFs. When compared with the “Conven (HWL)” scheme, the proposed scheme exhibits comparable performance but significantly less elapsed time per iteration. These results demonstrate that employing the fRIS can compensate for performance losses caused by reducing RIS elements, and alleviate the complexity of high-dimensional optimization.
@@
For the sensing performance evaluation, we demonstrate the fRIS beampattern under different weighting coefficients ($\alpha=0.1,0.5,0.9$), and compare with the ideal case and the reference beampattern in Figs. 3-3. It is observed that beams are consistently formed in the target directions. In the case of $\alpha=0.1$ where the system is biased towards communication, the beam is not as similar as the sensing-only case and exhibits higher sidelobe levels. In contrast, when $\alpha=0.5,0.9$ where the system is increasingly biased towards the sensing, the formed beams gradually align with the reference beampattern. While a discrepancy in beampattern energy exists between the “Proposed” and “Sensing-only” cases, it arises from the joint optimization process, where the inherent tradeoff introduces a mismatch between the covariance of the actual signal $\mathbf{v}$ and that of the reference signal $\mathbf{s}_{r}$. Consequently, the realized signal may differ from $\mathbf{s}_{r}$ in both shape and power. In addition, since $\mathbf{s}_{r}$ is a virtual design target and not physically transmitted, this mismatch does not violate energy conservation. On the other hand, Fig. 3 compares the beampattern of different approaches. It is observed that all the beampatterns have formed peaks in the target directions. Meanwhile, the proposed method achieves lower sidelobe levels compared to the conventional case and higher peaks compared to the “DPS” cases.
To further evaluate the sidelobe levels and the energy leakage, we demonstrate integrated sidelobe to mainlobe ratio (ISMR) [4] in Table III. Specifically, ISMR is computed as the ratio of the integrated power in the sidelobe region to that in the mainlobe region. Given the target directions $\left[-20^{\circ},5^{\circ},30^{\circ}\right]$, we set the mainlobe regions as $\left[-30^{\circ},-10^{\circ}\right]\cup\left[-5^{\circ},15^{\circ}\right]\cup\left[20^{\circ},40^{\circ}\right]$. It is observed that higher $\alpha$ yields lower ISMR, demonstrating better sidelobe suppression, while the proposed fRIS-aided design consistently achieves lower ISMR than the conventional RIS scheme with the same amount of element. For further sidelobe suppression, we may consider incorporating ISMR into optimization. Such enhancement is non-trivial and worth exploring in future work.
Under closely spaced targets, we simulate the beampatterns when two targets are placed at $0^{\circ}$ and another at $5^{\circ}$ or $10^{\circ}$. The aperture size of fRIS is set as $A\!\in\!\{8\lambda,12\lambda\}$ and the number of elements is set as $N\!\in\!\{16,64\}$. Fig. 4 reveals that larger angular separation improves beam distinguishability, while increasing aperture $A$ narrows the mainlobes. For instance, while $A\!=\!8\lambda$ leads to merged beams at $5^{\circ}$ separation, $A\!=\!12\lambda$ achieves separated peaks. Additionally, increasing $N$ can suppress the sidelobe and enhance the mainlobe gain.
TABLE III: ISMR under different settings.
$\alpha=0.1$ $\alpha=0.5$ (Default) $\alpha=0.9$
ISMR (dB) 1.49 -0.22 -0.51
Conven (SameNum): Conven (HWL): DPS
$N=25$ $N=81$
ISMR (dB) 2.65 -0.58 -0.18
(a)                                     (b)
@@
Figure 5: (a) Average BER (dB) versus different receive noise power (dBm) under different modulations and weighting coefficients. (b) Average BER (dB) versus different receive noise power (dBm) with different approaches.
@@
In addition, we evaluate the communication bit error rate (BER) of the system under different settings. Fig. 5(a) depicts the BER of the proposed scheme versus various noise power levels. As expected, the average BER decreases when the noise power decreases, and QPSK modulation exhibits lower BER when compared to higher-order modulations. Moreover, when $\alpha$ decreases, the BER improves as the system becomes more communication-biased. On the other hand, Fig. 5(b) compares the average BER of the proposed method to the benchmarks. As shown, all cases exhibit a decreasing BER trend as noise power level reduces, while the proposed method consistently achieves a lower BER compared to the baselines. In Fig. 6, the impact of the fRIS size on BER is evaluated, where we find that the larger fRIS aperture size yields lower BER. This is because increasing $A$ can form narrower beams, which increase the received energy for users.
Figure 6: Average BER (dB) versus different receive noise power (dBm) under different fRIS sizes.